OpenStudy (anonymous):

Integrate following integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}e^x e^{2x} e^{3x}...e^{200x}dx\]

OpenStudy (anonymous):

it would just be x,2x,3x....200x

OpenStudy (anonymous):

all have the +C part :)

OpenStudy (anonymous):

Damn I always forget those..haha

OpenStudy (zarkon):

\[e^x e^{2x} e^{3x}...e^{200x}=e^{x+2x+3x+\cdots+200x}\] \[=e^{x(1+2+3+\cdots+200)}\]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

Nope :(

OpenStudy (anonymous):

Try again guys :)

OpenStudy (anonymous):

you have some kind of choices?

OpenStudy (anonymous):

I'll give a hint, it requires a special series :)

myininaya (myininaya):

\[\frac{1}{\sum_{i=1}^{200}i}e^{(\sum_{i=1}^{200}i)x}+C\]

OpenStudy (anonymous):

law of exponential?

OpenStudy (zarkon):

just use the equality I gave above

OpenStudy (zarkon):

or myininaya answer ;)

OpenStudy (anonymous):

Close myininaya :)

OpenStudy (zarkon):

the sum is 20100

myininaya (myininaya):

\[\sum_{i=1}^{200}i=\frac{200(200+1)}{2}=100(200+1)=20000+100=20100\]

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

I devised that problem myself, sorry it wasn't a challenge myininaya :(

myininaya (myininaya):

lol